OpenStudy (anonymous):

A car cost $20,000 when it was purchased. The value of the car decreases by 8% each year. Find the rate of decay each month and select the correct answer below.

jimthompson5910 (jim_thompson5910):

Monthly Rate of Decay = (Annual Rate of Decay)/12

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@nincompoop @pooja195 @Astrophysics

OpenStudy (anonymous):

−0.006924% −0.006667% −0.666667% −0.6924% These are the answer choices..... I tried doing it, but my answer was not one of these

jimthompson5910 (jim_thompson5910):

Annual Rate of Decay = 8% Monthly Rate of Decay = (Annual Rate of Decay)/12 Monthly Rate of Decay = (8%)/12 Monthly Rate of Decay = ???

OpenStudy (anonymous):

My textbook confused me a lot sooooo ill give it my best shot, \[A=P(1+r ^{1/c})^{c*t}\] Is that the equation? @jim_thompson5910 If so, i just need help finding A, which i am very confused about

OpenStudy (anonymous):

@legends

jimthompson5910 (jim_thompson5910):

you're overthinking

jimthompson5910 (jim_thompson5910):

you just divide 8 by 12

OpenStudy (anonymous):

yes i know for the monthly rate btw wouldn't it btw 0.08/12 because it is 8% @jim_thompson5910

OpenStudy (anonymous):

oh wait, do you mean dividing 0.08/12 is all i do for the monthly rate

jimthompson5910 (jim_thompson5910):

well they leave things in percentages

jimthompson5910 (jim_thompson5910):

so (8%)/12 = 0.667%

OpenStudy (anonymous):

Oh ok thanks :)

jimthompson5910 (jim_thompson5910):

np